40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 130     Entry time: Tue May 25 17:01:57 2010
Author: tarac 
Type: Laser 
Category: Laser 
Subject: PMC cavity is locked 

      After adding another 1/2 plate to have P wave into the cavity, I can lock the PMC cavity. It's been 30 minutes so far.

There is only one transmitted beam now (there were two when I used S wave.) 

 

     When I work or knock on the table, sometime the beam switches to another mode (might be it's side band.)

It's very close to the main TEM00 mode. I need to adjust the DC offset a little bit to get back.

 The incoming power is 9 mW, and the Transmitted beam has ~ 6 mW.

Current setting:

RF Amp: 7V

Phase shift: 2.87 + 180

Gain: 27.91 dB

DC offset: -2.3 V

   

  

     I tried to pre align the ref cavity. It's harder than I thought, can't see the beam that well. I'll have to check the manual for the laser controller, so that I can

scan the RefCav when I align the beam into the cavity.

ELOG V3.1.3-